OpenStudy (osanseviero):

If I have 5 letters (A,B,C,D,E), and I want to form groups of three letters. a) How many can I form? (AAA, AAB, AAC, etc) b) How many without repeating? (ABC, ABD, etc)

OpenStudy (anonymous):

a) thats a combination of 3 from 5..or 5 C 3

OpenStudy (osanseviero):

So 125 in the first one, and 20 in the second one. Why is it n(n-1)?

OpenStudy (osanseviero):

:/

hero (hero):

If I'm not mistaken the first one is a permutation since elements are allowed to repeat. Which means the total number of three letter formations should be \(n^r=5^3 = 125\) The second question is a combination since elements are not allowed to repeat. So I imagine that it is \(5C3\)

OpenStudy (osanseviero):

Ok, let me check :)

OpenStudy (osanseviero):

I searched an in a permutation elements cant be repeated :/

hero (hero):

You should do more research on permutations.

OpenStudy (osanseviero):

So the basic formulas are \[n ^{r}\] if they can be repeated and NCr, when they cant be repeated?

hero (hero):

This resource might help clear things up a bit more: http://www.mathsisfun.com/combinatorics/combinations-permutations.html

OpenStudy (osanseviero):

Thanks :)

hero (hero):

yw

OpenStudy (osanseviero):

Yep, this helped me a lot :)

OpenStudy (osanseviero):

I am seeing that the formula of the combinations without repetition and permutation without repetition are the same. How can I know if this if a permutation or a combination? (the order matters here?)

hero (hero):

They're not the same :)

OpenStudy (osanseviero):

oops, not the same, I reduce by r one more time. But how can I know in this case if it is permutation or combination?

OpenStudy (osanseviero):

(from the question?, how can I know if order matters?

OpenStudy (osanseviero):

So it is 125 and 10 :)